OpenStudy (anonymous):

Does this polynomial have -3 and 2 as zeros? p(x)=2x^2+2x-12

jimthompson5910 (jim_thompson5910):

plug in x = -3 and tell me what you get

OpenStudy (anonymous):

6.66

jimthompson5910 (jim_thompson5910):

p(x)=2x^2+2x-12 p(-3)=2(-3)^2+2(-3)-12 p(-3)=2(9)+2(-3)-12 p(-3)=18-6-12 p(-3)=0

jimthompson5910 (jim_thompson5910):

So that proves that x = -3 is one root or zero

OpenStudy (anonymous):

What about p(x)=x^2-x-6?

jimthompson5910 (jim_thompson5910):

same idea: plug in the possible roots one at a time if you get 0, then it's confirmed to be a root

OpenStudy (anonymous):

I got 6...

jimthompson5910 (jim_thompson5910):

what are the potential roots you want to test?

OpenStudy (anonymous):

The same ones

OpenStudy (anonymous):

-3 & 2

jimthompson5910 (jim_thompson5910):

ok

jimthompson5910 (jim_thompson5910):

testing x = -3 p(x)=x^2-x-6 p(-3)=(-3)^2-(-3)-6 p(-3)=(9)-(-3)-6 p(-3)=9+3-6 p(-3)=6

jimthompson5910 (jim_thompson5910):

so x = -3 is NOT a root of p(x)=x^2-x-6

jimthompson5910 (jim_thompson5910):

testing x = 2 p(x)=x^2-x-6 p(2)=(2)^2-2-6 p(2)=4-2-6 p(2)=-4 and neither is x = 2

OpenStudy (anonymous):

Ok, thank you for explaining it! :)

jimthompson5910 (jim_thompson5910):

yw